# Evaluating Cooling-Center Coverage Using Persistent Homology of a Filtered Witness Complex

### v2

@@ Abstract.
Program in Applied & Computational Mathematics, Princeton University, Princeton, NJ (erinoneil@princeton.edu).
Department of Mathematics and Computer Science, College of the Holy Cross, Worcester, MA (stymochko@holycross.edu).
1. Introduction
@@
In a study that aimed to identify such protective measures, Bouchama found that visiting air conditioned environments was strongly associated with better health outcomes [5]. In fact, the meta-analysis showed that those who visited air conditioned places had an approximately 66% lower likelihood of heat-related mortality. However, approximately 28.6% of housing units in the United States do not have air conditioning [28]. Furthermore, certain populations are more likely to lack air conditioning due to limited financial and social resources [28]. Therefore, it is important for cities to offer free, accessible, public cooling centers, such as libraries, community centers, and senior centers [28] during extreme heat events. It is imperative not only to increase the quantity of cooling centers but also to strategically position them in areas of highest necessity, taking into account factors such as the demographics within the city and existing cooling center locations. Recent work shows that the distribution of cooling centers is not always optimized to maximize access [28].
1.1. Related Work and Contributions
@@
A tool from topological data analysis (TDA), persistent homology (PH), provides a complementary perspective to these existing techniques. Firstly, TDA can evaluate coverage across a range of parameter values which bypasses the arbitrary selection of a cut-off distance. Secondly, many traditional methods, such as those that rely solely on the information provided by an HVI, treat each census block as an isolated “island” when assessing the risk of its residents. In contrast, PH is able to incorporate spatial information by considering cooling centers in neighboring census blocks. Lastly, our methodology considers cooling centers not only within city limits, but also in neighboring cities. [^1] See Appendix A.2 for a detailed explanation of the data and how it was collected. In reality, individuals who live on the boundary of a given city may reside closer to cooling centers in neighboring cities than in their own. We therefore allow for the assumption that in an extreme heat event, individuals may visit neighboring cities to find shelter.
Persistent homology has shown success in many applications to spatial data [15, 24, 16, 9, 11]. Our method is inspired by that of Hickok et al. [23] which uses TDA to infer holes in voting site coverage. However, we use a different construction of the mathematical representation of space.
1.2. Organization of this paper
In Section 2 we introduce the relevant background from TDA and introduce on persistent homology. In Section 3, we present our persistent homology-based method as well as the HVI we use for comparison. We present and analyze the results in Section 4 and conclude with discussions of implications and limitations of our work in Section 5.
2. Background
@@ 2.1. Definitions
(1) $\mathcal{K}_{\alpha_{0}}\subseteq\mathcal{K}_{\alpha_{1}}\subseteq\cdots\subseteq\mathcal{K}_{\alpha_{n}}\,$,
where $\alpha_{0}<\alpha_{1}<\cdots<\alpha_{N}$ and $\mathcal{K}_{n}=\{\sigma\in\mathcal{K}:f(\sigma)<\alpha_{n}\}$. Equation 1 defines a filtration. We call $\alpha$ the filtration parameter.
2.1.1. Constructing a Simplicial Complex from Data
@@
The Čech complex is very useful in theory, but in practice is very computationally prohibitive. It is common to instead use the Vietoris–Rips (VR) complex as it is an approximation of the Čech complex and is significantly faster to compute. [^2] Computing the Čech complex requires knowledge of the metric space in order to check if the intersection of $k$ balls is nonempty. In comparison the VR complex requires only the pairwise distances between all points. The VR complex at filtration parameter $\alpha$ is the simplicial complex with vertices $X$ and simplices $\{x_{i_{0}},\ldots x_{i_{k}}\}$ if $d(x_{i_{j}},x_{i_{\ell}})<2\alpha$ for all $j$ and $\ell$. Just as with the Čech complex, the VR complex gives rise to a filtration when computed for an increasing sequence of filtration parameter values.
Here we use a different approach for creating a simplicial complex from data. In particular, we use the witness complex $\mathcal{W}$ because it allows for the incorporation of two sets of points, the landmarks $L$ and the witnesses $W$ [38]. Often the landmarks are chosen as a subset of the witnesses, however, this is not a requirement. [^3] The VR complex remains computationally prohibitive for large point sets. The witness complex can capture similar topological structure in a way that is more computationally tractable. In this setting, the full point cloud serves as the set of witnesses while a subset of the points serve as the landmarks. The resulting complex has fewer simplices than a Vietoris–Rips complex, thus speeding up computations. This, however, is not our motivation for using the witness complex. In our case, the landmarks and witnesses will come from two separate sets of points; the landmarks represent neighborhoods throughout a city and the witnesses represent the locations of the cooling centers. The landmarks form the vertices of the simplicial complex and witnesses are used to determine which higher dimensional simplices are present. That is, a $k$-simplex for $k>0$ is in the complex if it is “witnessed” by a point in $W$. This “witness” relation can be defined in different ways to create variants such as a weak or strong witness complex [38, 1, 36]. We define a filtered witness complex as a witness complex paired with a filter function $f:\mathcal{W}\to\mathbb{R}$ where $f(\sigma)$ is the distance from $\sigma$ to the closest witness. This gives rise to a filtration using the same definition as in Eqn. 1. We define our construction of the witness complex in Section 3.1.
2.2. Persistent Homology
@@
The birth simplex is the first simplex to appear in the filtration that creates a homology class, while the death simplex is the last simplex added that kills the homology class. Note that in general, a death simplex may not be unique; for example, if three vertices are all born at time 0 and the three pairwise edges are all added at the same filtration value, all three edges could be considered the death simplex. In real datasets, this is rarely an issue: distances are typically distinct, so edges and higher-dimensional simplices enter at unique filtration values.
@@
Figure 1. An example of $1$ census tract centroids (landmarks) and 7 cooling centers (witnesses). Note: this image was created based on a small subset of Boston, MA census tracts and the witness locations were generated randomly. To see the full city of Boston with the true locations of cooling centers, see Figure 11.
3. Methods
@@
3.1. Our Construction of the Witness Complex
The witness complex is a simplicial complex constructed based on two sets of points: landmarks, $L=\{\ell_{i}\}$, and witnesses, $W=\{w_{j}\}$. We will construct it as follows:
Definition 3.1.
Let $L$ be the set of landmark points, $W$ be the set of witness points, and $\alpha\in\mathbb{R}^{\geq 0}$ be a scale parameter. We define the witness complex constructed from $L$ and $W$ at scale $\alpha$ as follows:
@@ 3.1. Our Construction of the Witness Complex
Intuitively, this is a simplicial complex constructed with landmarks as the vertices (all of which are born at $\alpha=0$), and simplices built from the sets of vertices contained within distance $\alpha$ from the same witness. Note that this construction is a clique complex; that is, if all faces of a simplex are in the complex, the face is included as well. This makes computing the witness complex similar in computation time to the VR complex since all that is needed are the pairwise distances between all pairs of landmark and witness points. Given an increasing sequence of $\alpha$ values
@@
Figure 2. Example of select steps in the filtration. As shown in Figure 1, light green points represent the centroids (landmarks) of the geographic region and dark green diamonds (witnesses) represent the cooling center locations. All vertices (landmarks) appear in (a). (b)-(d) show the simplicial complex (in black) at various stages of the filtration. Edges are drawn between vertices when they are within a distance $\alpha$ of the same witness. Green discs show the radii at that filtration value. Note that the witnesses are drawn for easier visualization but they are not vertices in the simplicial complex.
@@ 3.1.2. Distance Computation
Figure 3. Heat vulnerability index (HVI) score maps for the four regions of interest.
3.2. Heat Vulnerability Index (HVI)
@@
The HVI is based on characteristics of the demographics and environment that have been shown to exacerbate the likelihood that a person will suffer negative health consequences from an extreme heat event. Since the demographic data is not available at the census block level, we define this score on census tracts. The characteristics of a census tract $T$ that we use are the typical afternoon temperature $F(T)$ in Fahrenheit, [^8] Note that afternoon temperature data was collected at 3pm in the respective city using a combination of Sentinel-2 satellite data and sensors mounted on the bikes and cars of volunteers [31] the percentage of the area not covered by tree canopy $C(T)$, and the number of younger and older residents, $Y(T)$ and $S(T)$ respectively. [^9] Note that the National Integrated Heat Health Information System defines “younger” to mean under 5 years old and “older” to mean over 65 years old [31]. Because these variables have different units of measurement, we standardize them before creating a composite score. The HVI $HVI(T)$ of census tract $T$ is
(2) $HVI(T)=\frac{F(T)-\mu_{F}}{\sigma_{F}}+\frac{C(T)-\mu_{C}}{\sigma_{C}}+\frac{Y(T)-\mu_{Y}}{\sigma_{Y}}+\frac{S(T)-\mu_{S}}{\sigma_{S}}\,$,
@@ 3.2.1. Evaluating Multicollinearity
Cutter et al. explains that an essential component of building a HVI involves testing for multicollinearity among demographic variables [10]. We examine the four variables used in this study and test for independence using variance inflation factor (VIF) [22]. The full VIF results can be found in the Appendix in Table 4, however we briefly discuss them here. The VIF analysis for Austin, Miami, and Boston indicate negligible multicollinearity for all four variables. The results for Portland indicate the afternoon temperature and tree canopy have non-negligible multicollinearity and thus are likely not independent of each other. Because of this, the HVI results for Portland should be considered with caution. We further analyze the independence and multicollinearity using pairwise correlations in Appendix A.3 and Figures 1619.
@@
Figure 4. Persistence diagram for our four geographic locations. The birth and death times are in units of kilometers. Note that the axes are not on the same scale.
3.2.2. Limitations
@@
While the purpose of the HVI is to serve as a comparison to our topological approach, we will briefly discuss its limitations. We limit our HVI to only include four variables for the sake of simplicity, however, the U.S. Urban Heat Island Mapping Campaign dataset contains sever other variables that could be incorporated. This includes features such as average morning and evening temperatures, the number of individuals living in poverty, the median income, and the percentage of the population that is a minority [^10] Defined by the U.S. Urban Heat Island Mapping Campaign to mean non-white.. Because we chose two demographic variables, it is possible the HVI results are biased towards places with a higher population. For example, places that are cooler but have a higher population of individuals <5<5 years old or >65>65 years old may have a higher HVI than a place that is warmer but has a lower population of vulnerable individuals. Beyond which variables are included, there are modifications that could be made to the approach. While the unweighted $z$-score approach is standard in the existing literature, it weights each variable equally and it increases linearly with the increase of any one variable. Because measuring heat vulnerability is multifaceted, it is unclear if these are valid assumptions. [^11] One can explore the HVI and incorporate different weightings and combinations of variables using code in our Github repository.
@@
Figure 5. Box plots for each of the four geographic locations of interest of (a) death times of the 0D, (b) death times of 1D homology classes, and (c) the lifetimes of the 1D homology classes. A tables of the corresponding means and standard deviations can be found in Table 1.
Table 1. Mean and standard deviation (SD) of: dimension 0 death times, dimension 1 death times, and dimension 1 lifetimes.
0D Death Time 1D Death Time 1D Lifetime
City Mean SD Mean SD Mean SD
Austin 1.939 1.428 2.243 0.718 0.060 0.041
Boston 0.745 0.454 1.08 0.554 0.144 0.172
Miami 1.15 0.518 1.345 0.609 0.167 0.160
Portland 1.418 0.796 2.425 0.631 0.258 0.343
4. Results
@@
We test our methodology on four locations in the United States that offer unique climates and demographics. Those locations include central Boston, MA; central Austin, TX; Portland, OR; and Miami, FL. These cities differ in various ways, for example Miami and Boston are on the water, while Austin is landlocked. Additional differences include humidity, which can increase how warm a temperature feels, and the populations of vulnerable demographics such as older adults. It is important to note that all the selected cities are vulnerable to the urban heat island effect [31].
4.1. HVI
@@
Based on the HVI maps in Figure 3, we can identify the regions of higher HVI which indicate regions more vulnerable to heat related mortality. In particular, we can see individual census tracts in central Austin, northern Boston, western Miami, and eastern Portland with particularly high HVIs. We can see based on the box plots of the HVI values in Figure 6 that all cities have similar medians, although the HVI values in Miami are slightly higher. A table of statistics of the HVI values can be found in Appendix A.3 in Table 5. Additionally, Figures 1215 show heatmaps of each of the individual variables used in the HVI computation. These can be used to identify regions where one particular indictor of vulnerability may be high even if the composite HVI is not.
@@
Figure 6. Boxpot of HVI across all census tracts within a given city.
4.2. Topological Approach

### v3

@@ Abstract
1 Introduction
@@
In a study that aimed to identify such protective measures, Bouchama et al. found that visiting air conditioned environments was strongly associated with better health outcomes [5]. In fact, the meta-analysis showed that those who visited air conditioned places had an approximately 66% lower likelihood of heat-related mortality. However, approximately 28.6% of housing units in the United States do not have air conditioning [29]. Furthermore, certain populations are more likely to lack air conditioning due to limited financial and social resources [29]. Therefore, it is important for cities to offer free, accessible, public cooling centers, such as libraries, community centers, and senior centers [29] during extreme heat events. It is imperative not only to increase the quantity of cooling centers but also to strategically position them in areas of highest necessity, taking into account factors such as the demographics within the city and existing cooling center locations. Recent work shows that the distribution of cooling centers is not always optimized to maximize access [29].
1.1 Related Work and Contributions
@@
A tool from topological data analysis (TDA), persistent homology (PH), provides a complementary perspective to these existing techniques. Firstly, TDA can evaluate coverage across a range of parameter values which bypasses the arbitrary selection of a cut-off distance. Secondly, many traditional methods, such as those that rely solely on the information provided by an HVI, treat each census block as an isolated “island” when assessing the risk of its residents. In contrast, PH is able to incorporate spatial information by considering cooling centers in neighboring census blocks. Lastly, our methodology considers cooling centers not only within city limits, but also in neighboring cities. [^1] See Appendix A.2 for a detailed explanation of the data and how it was collected. In reality, individuals who live on the boundary of a given city may reside closer to cooling centers in neighboring cities than in their own. We therefore allow for the assumption that in an extreme heat event, individuals may visit neighboring cities to find relief.
Our method is inspired by that of Hickok et al. [24] which uses TDA to infer holes in voting site coverage. However, we use a different construction of the mathematical representation of space. Other researchers have also addressed questions of resource coverage using PH. In particular, Gonzalez-Cruz et al. [21] used PH to quantify the accessibility of Planned Parenthood and other health care clinics in California to identify regions lacking adequate health care facilities or those vulnerable in the case of clinics being defunded. Persistent homology has shown success in many applications to spatial data beyond problems in resource coverage [15, 25, 16, 9, 11].
1.2 Organization of this paper
In Section 2 we introduce the relevant background from TDA; specifically, we introduce persistent homology. In Section 3, we present our persistent homology-based method as well as the HVI we use for comparison. We present and analyze the results in Section 4 and conclude with discussions of implications and limitations of our work in Section 5.
2 Background
@@ 2.1 Definitions
(2.1) $\mathcal{K}_{\alpha_{0}}\subseteq\mathcal{K}_{\alpha_{1}}\subseteq\cdots\subseteq\mathcal{K}_{\alpha_{n}}\,$,
where $\alpha_{0}<\alpha_{1}<\cdots<\alpha_{N}$ and $\mathcal{K}_{n}=\{\sigma\in\mathcal{K}:f(\sigma)<\alpha_{n}\}$. Equation 2.1 defines a filtration. We call $\alpha$ the filtration parameter.
2.1.1 Constructing a Simplicial Complex from Data
@@
The Čech complex is very useful in theory, but in practice is very computationally prohibitive. It is common to instead use the Vietoris–Rips (VR) complex as it is an approximation of the Čech complex and is notably faster to compute. [^2] Computing the Čech complex requires knowledge of the metric space in order to check if the intersection of $k$ balls is nonempty. In comparison, the VR complex requires only the pairwise distances between all points. The VR complex at filtration parameter $\alpha$ is the simplicial complex with vertices $X$ and simplices $\{x_{i_{0}},\ldots x_{i_{k}}\}$ if $d(x_{i_{j}},x_{i_{\ell}})<2\alpha$ for all $j$ and $\ell$. Just as with the Čech complex, the VR complex gives rise to a filtration when computed for an increasing sequence of filtration parameter values.
Here we use a different approach for creating a simplicial complex from data. In particular, we use the witness complex $\mathcal{W}$ because it allows for the incorporation of two sets of points, the landmarks $L$ and the witnesses $W$ [40]. Often the landmarks are chosen as a subset of the witnesses, however, this is not a requirement. [^3] The VR complex remains computationally prohibitive for large point sets. The witness complex can capture similar topological structure in a way that is more computationally tractable. In this setting, the full point cloud serves as the set of witnesses while a subset of the points serve as the landmarks. The resulting complex has fewer simplices than a Vietoris–Rips complex, thus speeding up computations. This, however, is not our motivation for using the witness complex. In our case, the landmarks and witnesses will come from two separate sets of points; the landmarks represent neighborhoods throughout a city and the witnesses represent the locations of the cooling centers. The landmarks form the vertices of the simplicial complex and witnesses are used to determine which higher dimensional simplices are present. That is, a $k$-simplex for $k>0$ is in the complex if it is “witnessed” by a point in $W$. This “witness” relation can be defined in different ways to create variants such as a weak or strong witness complex [40, 1, 38]. We define a filtered witness complex as a witness complex paired with a filter function $f:\mathcal{W}\to\mathbb{R}$ where $f(\sigma)$ is the distance from $\sigma$ to the closest witness. This gives rise to a filtration using the same definition as in Eqn. 2.1. We define our construction of the witness complex in Section 3.1.
2.2 Persistent Homology
@@
The birth simplex is the first simplex to appear in the filtration that creates a homology class, while the death simplex is the last simplex added that kills the homology class. In general, a death simplex may not be unique; for example, if three vertices are all born at time 0 and the three pairwise edges are all added at the same filtration value, all three edges could be considered the death simplex. In real datasets, this is rarely an issue: distances are typically distinct, so edges and higher-dimensional simplices enter at unique filtration values. Another concern with death simplices is that they are not stable; that is, small perturbations in the data can cause the death simplex to change locations.
@@
Figure 1: An example of $1$ census tract centroids (landmarks) and 7 cooling centers (witnesses). Note: this image was created based on a small subset of Boston, MA census tracts and the witness locations were generated randomly. To see the full city of Boston with the true locations of cooling centers, see Figure 11.
3 Methods
@@
3.1 Our Construction of the Witness Complex
The witness complex is a simplicial complex constructed based on two sets of points: landmarks, $L=\{\ell_{i}\}$, and witnesses, $W=\{w_{j}\}$. We will construct it as follows: {definition*} Let $L$ be the set of landmark points, $W$ be the set of witness points, and $\alpha\in\mathbb{R}^{\geq 0}$ be a scale parameter. We define the witness complex constructed from $L$ and $W$ at scale $\alpha$ as follows:
@@ 3.1 Our Construction of the Witness Complex
Intuitively, this is a simplicial complex constructed with landmarks as the vertices (all of which are born at $\alpha=0$), and simplices built from the sets of vertices contained within distance $\alpha$ from the same witness. Given an increasing sequence of $\alpha$ values
@@
Figure 2: Example of select steps in the filtration. As shown in Figure 1, light green points represent the centroids (landmarks) of the geographic region and dark green diamonds (witnesses) represent the cooling center locations. All vertices (landmarks) appear in (a). (b)-(d) show the simplicial complex (in black) at various stages of the filtration. Edges are drawn between vertices when they are within a distance $\alpha$ of the same witness. Green discs show the radii at that filtration value. Note that the witnesses are drawn for easier visualization but they are not vertices in the simplicial complex.
@@ 3.1.2 Distance Computation
Figure 3: Heat vulnerability index (HVI) score maps for the four regions of interest.
3.2 Heat Vulnerability Index (HVI)
@@
The HVI is based on characteristics of the demographics and environment that have been shown to exacerbate the likelihood that a person will suffer negative health consequences from an extreme heat event. Since the demographic data is not available at the census block level, we define this score on census tracts. The characteristics of a census tract $T$ that we use are the typical afternoon temperature $F(T)$ in Fahrenheit, [^8] Note that afternoon temperature data was collected at 3pm in the respective city using a combination of Sentinel-2 satellite data and sensors mounted on the bikes and cars of volunteers [33] the percentage of the area not covered by tree canopy $C(T)$, and the number of younger and older residents, $Y(T)$ and $S(T)$ respectively. [^9] Note that the National Integrated Heat Health Information System defines “younger” to mean under 5 years old and “older” to mean over 65 years old [33]. Because these variables have different units of measurement, we standardize them before creating a composite score. The HVI of census tract $T$ is
(3.1) $HVI(T)=\frac{F(T)-\mu_{F}}{\sigma_{F}}+\frac{C(T)-\mu_{C}}{\sigma_{C}}+\frac{Y(T)-\mu_{Y}}{\sigma_{Y}}+\frac{S(T)-\mu_{S}}{\sigma_{S}}\,$,
@@ 3.2.1 Evaluating Multicollinearity
Cutter et al. explains that an essential component of building a HVI involves testing for multicollinearity among demographic variables [10]. We examine the four variables used in this study and test for independence using variance inflation factor (VIF) [23]. The full VIF results can be found in the Appendix in Table 4, however we briefly discuss them here. The VIF analysis for Austin and Miami indicate that the low amount of multicollinearity is acceptable for the proposed analysis for all four variables. The results for Portland indicate the afternoon temperature and tree canopy have non-negligible multicollinearity and thus are likely not independent of each other. Because of this, the HVI results for Portland should be considered with caution. We further analyze the independence and multicollinearity using pairwise correlations in Appendix B and Figures 13 and 14. This analysis reveals a strong pairwise correlation between the afternoon temperature and tree canopy in Boston. Thus, the HVI results for Boston must also be considered with caution.
@@
Figure 4: Persistence diagram for our four geographic locations. The birth and death times are in units of kilometers. Note that the axes are not on the same scale.
3.2.2 Limitations
@@
While the purpose of the HVI is to serve as a comparison to our topological approach, we will briefly discuss its limitations. We limit our HVI to only include four variables for the sake of simplicity, however, the U.S. Urban Heat Island Mapping Campaign dataset contains sever other variables that could be incorporated. This includes features such as average morning and evening temperatures, the number of individuals living in poverty, the median income, and the percentage of the population that is a minority [^10] Defined by the U.S. Urban Heat Island Mapping Campaign to mean non-white.. Because we chose two demographic variables, it is possible the HVI results are biased towards places with a higher population. For example, places that are cooler but have a higher population of individuals <5<5 years old or >65>65 years old may have a higher HVI than a place that is warmer but has a lower population of vulnerable individuals. Beyond which variables are included, there are modifications that could be made to the approach. While the unweighted $z$-score approach is standard in the existing literature, it weighs each variable equally and it increases linearly with the increase of any one variable. Because measuring heat vulnerability is multifaceted, it is unclear if these are valid assumptions. [^11] One can explore the HVI and incorporate different weightings and combinations of variables using code in our Github repository.
@@
Figure 5: Box plots for each of the four geographic locations of interest of (a) death times of the 0D homology classes, (b) death times of 1D homology classes, and (c) the lifetimes of the 1D homology classes. A table of the corresponding means and standard deviations can be found in Table 1.
4 Results
@@
We test our methodology on four locations in the United States that offer unique climates and demographics. Those locations include central Boston, MA; central Austin, TX; Portland, OR; and Miami, FL. These cities differ in various ways, for example Miami and Boston are on the water, while Austin is landlocked. Additional differences include humidity, which can increase how warm a temperature feels, and the populations of vulnerable demographics such as older adults. All of the selected cities are vulnerable to the urban heat island effect [33].
4.1 HVI
@@
Based on the HVI maps in Figure 3, we can identify the regions of higher HVI which indicate regions more vulnerable to heat related mortality. In particular, we can see individual census tracts in central Austin, northern Boston, western Miami, and eastern Portland with particularly high HVIs. We can see based on the box plots of the HVI values in Figure 6 that all cities have similar medians, although the HVI values in Miami are slightly higher. A table of statistics of the HVI values can be found in the last column of Table 1. Additionally, Figure 12 shows heatmaps of each of the individual variables used in the HVI computation for each city. These can be used to identify regions where one particular indictor of vulnerability may be high even if the composite HVI is not.
@@
Figure 6: Boxplot of the HVIs across all census tracts within a given city. A table of the corresponding means and standard deviations can be found in Table 1.
4.2 Topological Approach
Table 1: Mean and standard deviation (SD) of: dimension 0 death times, dimension 1 death times, dimension 1 lifetimes, and HVI. Note that the units of death times and lifetimes are kilometers. The HVI is unitless due to it being an unweighted z-socre.
0D Death 1D Death 1D Lifetime HVI
Time Times
City Mean $\pm$ SD Mean $\pm$ SD Mean $\pm$ SD Mean $\pm$ SD
Austin 1.939 $\pm$ 1.428 2.812 $\pm$ 0.820 0.152 $\pm$ 0.141 7.819 $\pm$ 1.746
Boston 0.745 $\pm$ 0.454 1.111 $\pm$ 0.567 0.146 $\pm$ 0.174 1.326 $\pm$ 1.499
Miami 1.149 $\pm$ 0.519 1.677 $\pm$ 0.532 0.217 $\pm$ 0.156 8.715 $\pm$ 2.051
Portland 1.418 $\pm$ 0.796 2.268 $\pm$ 0.530 0.259 $\pm$ 0.312 7.046 $\pm$ 1.805